# Learning in Structured Stackelberg Games

### v1

@@ Appendix B Details and Proofs from Section 4
By the boundedness of the utility function we have that the expectation above lies within $[-1,1]$. Therefore, we get:
@@
But from Theorem 6 in [17], we know that for any $\epsilon,\delta>0$, there exists constant $C_{2}$ such that for any $m\geq C_{2}\left(\frac{d_{N}(\pazocal{H}))\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$, we have that for any $h_{\text{ERM}}$ that satisfies equation (3):
@@
B.1 Proof of Theorem 4.12
See 4.12
@@
$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\pi_{\text{S-ERM}})\geq u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)$ (5)
where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$. In other words, it suffices to prove that all the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ with $u_{\mathscr{D}}(\pi_{h})<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon$ have low utility in the training sample too, that is, $u_{S}(\pi_{h})\leq u_{S}(\pi_{h^{*}})$.
@@
  • For $i$ from $1$ to $m$, flip a fair coin.
@@
By Claim B.2, $\mathbb{P}[B^{\prime\prime}]=\mathbb{P}[B^{\prime}]$ so, it suffices to upper bound $\mathbb{P}[B^{\prime\prime}]$. Notice that the first probability is over the choice of $S$, $S^{\prime}$, and the randomness of SwapR, while the second is over the choice of $S$ and $S^{\prime}$.
@@
By definition, we have that:
@@
Using the union bound, we get that:
@@
Combining all these, we get that $\mathbb{P}[B]\leq\delta$, whenever $2\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}\leq\delta$, which proves that:
@@
It remains to show that $\Pi_{\pazocal{H},\pazocal{G}}\leq d_{SN,\pazocal{G}}(\pazocal{H})\ln(K)$. We prove this using the Lemma 4.13, which states that $\Pi_{\pazocal{H},\pazocal{G}}[m]\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$. So we get that:
@@
Rearranging, we obtain the requirement that:
@@
Using lemma B.5, a sufficient condition for the above to hold is that:
@@
And a sufficient condition for the above is:
@@
The above holds if there exists a constant $C$, large enough such that:
@@
B.2 Proof of Lemma 4.13
See 4.13
@@
We will show that for $|Z|=m$:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ |$\leq|\{B\subseteq\pazocal{Z}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (6)
$\leq\sum_{i=1}^{d_{SN,\pazocal{G}}(\pazocal{H})}\binom{m}{i}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (7)
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (8)
Notice that equation (7) holds by enumerating all possible subsets of $\pazocal{Z}$ that can be SN-shattered by $\pazocal{H}$ up to $d_{SN,\pazocal{G}}(\pazocal{H})$, which is the largest size subset that $\pazocal{H}$ can shatter. So it suffices to prove equation (6). We proceed by induction on the size of $|\pazocal{Z}|$.
Base Case: For $|\pazocal{Z}|=1$, we have that $|\Pi_{\pazocal{H},\pazocal{G}}|\leq K$ so (6) holds.
Inductive Hypothesis: For $|\pazocal{Z}|=m$ assume that (6) holds.
Inductive Step: We will show (6) for $|S|=m+1$. Consider the set $\pazocal{Z}=\{\mathbf{z}_{0},\mathbf{z}_{1},\cdots,\mathbf{z}_{m}\}$ and we define the following notation:
$\Lambda^{(i)}=\{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}|\pi_{h}(\mathbf{z}_{0})=i\}$
$\Lambda^{(i,j)}=\{\pi_{h}\in\Lambda^{(i)}|\exists\pi_{h}^{\prime}\in\Lambda^{(j)}\;\text{s.t.}\;\pi_{h}(\mathbf{z})=\pi_{h}^{\prime}(\mathbf{z})\;\forall\mathbf{z}\in\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\}$
For $i,j\in[K]$ also denote $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}$ to be all the policies $\pi_{f}:\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\to[K]$ that can be extended to a policy in $\Pi_{\pazocal{H},\pazocal{G}}$ both by defining $\pi_{f}(\mathbf{z}_{0})=i$ and by defining $\pi_{f}(\mathbf{z}_{0})=j$. Then we claim the following:
$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|,$ (9)
where $|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$ denotes the restriction of $\Pi_{\pazocal{H},\pazocal{G}}$ to the set $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$.
(Proof of equation (9)). Consider the $\ell$’th labeling of elements in $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$ that is consistent with a policy in $\Pi_{\pazocal{H},\pazocal{G}}$. Let $\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}$ denote the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ that are consistent with that labeling. Then we have that:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ $=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}$
$=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}1+\left(\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}\right)-1$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1$
$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}-1+\sum_{j=1}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\left(\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1\right)$
$\hskip 85.00012pt+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|.$ (10)
We proceed with the induction. From the inductive hypothesis we have that:
$|\displaystyle|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})},$ (11)
@@
where the last inequality holds from equation (8). Also, notice that each $\Lambda^{(i,j)}$ SN-shatters $\mathbf{z}_{0}$ by definition. So $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}$ cannot shatter a set larger than $d_{SN,\pazocal{G}}(\pazocal{H})-1$. Therefore, by the inductive hypothesis we get that:
$|\displaystyle|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}\;\text{SN-shatters}\;B\}|\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$ (12)
Combining inequalities (10), (11), and (12), we get that:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ $\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
$=m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
$=m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}(m+1)$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$

### v2

@@ B.2.1 Vanilla ERM learner
By the boundedness of the utility function we have that the expectation above lies within [1,1][-1,1][ - 1 , 1 ]. Therefore, we get
@@
But from Theorem 6 in [22], we know that for any $\epsilon,\delta>0$, there exists constant $C_{2}$ such that for any $m\geq C_{2}\left(\frac{d_{N}(\pazocal{H})\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$, we have that for any $h_{\text{ERM}}$ that satisfies Equation˜3
@@
B.2.2 Stackelberg ERM learner
In this section, we show how to improve upon the sample complexity bound shown in Proposition˜B.13, by leveraging the interaction between the Stackelberg game and the hypothesis class. Consider the following hypothesis class of possible leader policies
$\Pi_{\pazocal{H},\pazocal{G}}=\Bigl{\{}\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})|h\in\pazocal{H},\pi_{h}(\mathbf{z})\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))\Bigl{\}}$
and the following variation of an ERM rule.
Definition B.14 (Stackelberg ERM).
Consider a hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ and a Stackelberg game, $\pazocal{G}$. We define the Stackelberg ERM learning algorithm, $\textsc{Alg}_{\text{S-ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times[K])^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, outputs the policy $\pi_{\text{S-ERM}}:\pazocal{Z}^{m}\to\Delta(\pazocal{A})$ that satisfies
$\pi_{\text{S-ERM}}\in\arg\max_{\pi\in\Pi_{\pazocal{H},\pazocal{G}}}\frac{1}{m}\sum_{i=1}^{m}u(\mathbf{z}_{i},\pi(\mathbf{z}_{i}),b_{f_{i}}(\mathbf{z}_{i},\pi(\mathbf{z}_{i})).$ (4)
Definition B.15 (Stackelberg-Natarajan Dimension).
Consider a hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$, a Stackelberg game, $\pazocal{G}$, and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ SN-shatters $S$ with respect to game $\pazocal{G}$, if there exist $\phi_{1},\phi_{2}:S\to\Delta(\pazocal{A})$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that
Theorem B.16.
Proof.
Example B.17.
Consider the hypothesis class of conjunctions, where each hypothesis is defined over a set of Boolean variables, $\pazocal{Z}=\{z_{1},\cdots,z_{n}\}$, such that $\pazocal{H}_{\text{conj}}=\{h:\{0,1\}^{n}\to\{0,1\}|h(z)=\land_{i\in S}z_{i}\land_{j\in T}z_{j},S,T\subseteq[n],S\cap T=\emptyset\}$. Since each hypothesis outputs in binary, this hypothesis class could naturally be considered in a Stackelberg game with two follower types, $f^{(0)},f^{(1)}$. The Natarajan dimension in this case reduces to the well-studied notion of the VC-dimension, which is known to be equal to $n$ for $\pazocal{H}_{\text{conj}}$. At the same time, the $d_{SN,\pazocal{G}}(\pazocal{H}_{\text{conj}})$ will be upper bounded by the size of the set $\{z\in\{0,1\}^{n}|\pi_{*}^{(0)}(z)\neq\pi_{*}^{(1)}(z)\}$, which could be as small as 111 for two distinct follower types.
Theorem B.18.
Consider the Stackelberg ERM policy, $\pi_{\text{S-ERM}}$, as in Definition˜B.15 and let $h_{\text{S-ERM}}$ be the corresponding hypothesis in $\pazocal{H}$. Then universally for any such policy, there exists constant $C$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ and distribution $\mathscr{D}$,
Proof Sketch: The proof follows from a similar argument as that of the sample complexity bounds of multiclass learning [22]. The only part of the proof that does not apply here is Natarajan’s lemma. We prove an appropriate generalization in Lemma˜B.24.
@@ B.2.2 Stackelberg ERM learner
$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\pi_{\text{S-ERM}})\geq u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right).$ (5)
In other words, it suffices to prove that all the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ with $u_{\mathscr{D}}(\pi_{h})<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon$ have low utility in the training sample too, that is, $u_{S}(\pi_{h})<u_{S}(\pi_{h^{*}})$.
@@
  • For $i$ from 111 to $m$, flip a fair coin.
@@
By Lemma˜B.20, $\mathbb{P}[B^{\prime\prime}]=\mathbb{P}[B^{\prime}]$ so, it suffices to upper bound $\mathbb{P}[B^{\prime\prime}]$. Notice that the first probability is over the choice of $S$, $S^{\prime}$, and the randomness of SwapR, while the second is over the choice of $S$ and $S^{\prime}$.
@@
By definition, we have that
@@
Using the union bound, we get that
@@
Combining all these, we get that $\mathbb{P}[B]\leq\delta$, whenever $2\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}\leq\delta$, which proves that
@@
It remains to show that $\Pi_{\pazocal{H},\pazocal{G}}\leq d_{SN,\pazocal{G}}(\pazocal{H})\ln(K)$. We prove this using Lemma˜B.24, which states that $\Pi_{\pazocal{H},\pazocal{G}}[m]\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$. So we get that
@@
Rearranging, we obtain the requirement that
@@
Using Lemma˜B.23, a sufficient condition for the above to hold is that
@@
And a sufficient condition for the above is
@@
The above holds if there exists a constant $C$, large enough such that
@@
Lemma B.24.
For a given class $\pazocal{H}$ and Stackelberg game $\pazocal{G}$, let $d_{SN,\pazocal{G}}(\pazocal{H})=d$. Then for a set of contexts $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{|S|}\}$, let $\Pi_{\pazocal{H},\pazocal{G}}(S)=\{(\pi(\mathbf{z}_{1}),\dots,\pi(\mathbf{z}_{|S|}))\;\;|\;\;\pi\in\Pi_{\pazocal{H}},\pazocal{G}\}$. It holds that
$|\Pi_{\pazocal{H},\pazocal{G}}(S)|\leq|S|^{d}\cdot K^{2d}.$ (6)
@@
[…]
We will now show Equation˜6 for $|S|=m+1$. First we fix a set $S=\{\mathbf{z}_{0},\mathbf{z}_{1},\cdots,\mathbf{z}_{m}\}$. For simplicity and without loss of generality we write $\Pi_{\pazocal{H},\pazocal{G}}(S)$ as $\Pi_{\pazocal{H},\pazocal{G}}$. We now define the following
$\Lambda^{(i)}=\{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}|\pi_{h}(\mathbf{z}_{0})=\mathbf{x}_{*}^{(i)}\}$
$\Lambda^{(i,j)}=\{\pi_{h}\in\Lambda^{(i)}|\exists\pi_{h}^{\prime}\in\Lambda^{(j)}\;\text{s.t.}\;\pi_{h}(\mathbf{z})=\pi_{h}^{\prime}(\mathbf{z})\;\forall\mathbf{z}\in\pazocal{Z}\},$
where $\mathbf{x}_{*}^{(i)}$ is the leader’s optimal strategy against $f^{(i)}$, given context $\mathbf{z}_{0}$. We also define
$\Pi_{0}=\Pi_{\pazocal{H},\pazocal{G}}\setminus\left(\cup_{i\neq j}\Lambda^{(i,j)}\right).$
Then it immediately follows that
$|\Pi_{\pazocal{H},\pazocal{G}}(S)|\leq|\Pi_{0}(S)|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}(S)|.$ (7)
@@
By definition we know that each policy in $\Pi_{0}$, produces a distinct set of optimal strategies on the contexts in $S\setminus\{\mathbf{z}_{0}\}$. Therefore, $|\Pi_{0}(S)|=|\Pi_{0}(S\setminus\{\mathbf{z}_{0}\})|$. Also, $\Pi_{0}\subset\Pi_{\pazocal{H}},\pazocal{G}$ so $\Pi_{0}$ cannot SN-shatter a set of size greater than $d$. Therefore, from the inductive hypothesis we get that
$|\Pi_{0}(S)|=|\Pi_{0}(S\setminus\{\mathbf{z}_{0}\})|\leq m^{d}\cdot K^{2d}.$ (8)
Next, we argue that each $\Lambda^{(i,j)}$ cannot SN-shatter a set of size greater than $d-1$. To see this, we let $L=\{\tilde{\mathbf{z}}_{1},\dots,\tilde{\mathbf{z}}_{|}L|\}$ be a set that $\Lambda^{(i,j)}$ can SN-shatter. Then $L$ cannot include $\mathbf{z}_{0}$, since $\pi(\mathbf{z}_{0})=\mathbf{x}_{*}^{(i)}$ for all $\pi\in\Lambda^{(i,j)}$. However, $\Pi_{\pazocal{H},\pazocal{G}}$ can SN-shatter $L\cup\{\mathbf{z}_{0}\}$. So $|L|\leq d-1$. Therefore, by the inductive hypothesis we get
$|\Lambda^{(i,j)}(S)|\leq m^{d-1}\cdot K^{2(d-1)}$ (9)
Combining inequalities (7), (8), and (9), we get that
$|\Pi_{\pazocal{H},\pazocal{G}}(S)|$ $\leq m^{d}\cdot K^{2d}+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}m^{d-1}\cdot K^{2(d-1)}$
$\leq m^{d}\cdot K^{2d}+m^{d-1}\cdot K^{2d}$
$=m^{d-1}K^{2d}(m+1)$
$\leq(m+1)^{d}K^{2d}$